INAF LBC+LUCIFER queue observing, 2010 October 10-11

Observer: V. Testa, K.Boutsia, M. Fumana
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O.Kuhn
Instrument: LBC - LUCIFER

Summary

Clear sky with bad seeing at the beginning of the night
Time lost because of guiding problems with LBC
A MOS script might be compromised because of LUCIFER problems

Data Summary

Project Objects Filter ExpTime NotesSorted ascending
Cresci DSF2237a Ks 73min  
Cresci DSF2237b Ks 75min  
Turatto SN2009kr J, H, Ks 640 s + 896 s + 896 s  
Turatto SN2010al J, H, Ks 640 s + 896 s + 896 s  
Pentericci Cluster1 200H+K 1h data may be compromised
Phot_STD FS152 Ks   for the Cresci program
Clementini AndXXI - AndXIX B-bessel, V-bessel 21min + 21min seeing >1.3'' and guiding problems

Details

Take some LUCIFER Dark frames in the afternoon for tests
  • 2s (Dark_001)
  • 300s (Dark_002)
  • 60s (Dark_003)
  • 60s (Dark_004) - during switch just for test

Begin night with LBC

01:13 - Skyflat Test LBC - B(blue) - V(red)
  • 3sec B- 1 sec V saturated
  • 0.3x(3 e 1)sec B - V
  • 1.5 sec B - 0.5 sec V
  • PA 180 - 3sec B + 3sec V

01:32 - Slew to pointing star ACT0219
  • Pointing correction OK + mirror adjust OK

01:39 - slew to PHOT_STD - SA113 dofpia,
  • manually give Z7 = -2000(Red+Blue) and Z8=-1000(Red)
  • dofpia,/First
  • sky is still bright - too few pupils found
  • manually give Z4=+10000 (Red ) - Z4=+5000(Blue)

01:53 - Hit limits in red mirror - OB interrupted
  • coma free pointing adjustment on dx

01:57 - restart OB for STD - SA113_B_V

02:01 - Slew to second standard - SA110 _B_V * seeing ~1.3" * temperature fell abruptly 1.5deg

02:06 - Slew to target AndXXI (Clementini)
  • dofpia
  • seeing still bad (1.3"-1.5")
  • Blue converged after 4 iterations, red after 7 stopped with only Z4 slightly out of limits

02:20 - AndXXI OB1
  • sources are slightly elongated - seeing 1.6"(blue)-1.5"(red)
  • guiding problems?? - out of focus?? - just bad seeing??

02:45 - Slew to AndXIX - dofpia
  • stopped after 6 iterations - close to converge

02:55 - AndXIX OB1
  • images still elongated
  • guiding star (especially blue) is moving around a lot!
  • it is a guiding problem (blue is the master)

03:20 - OB finished but we have to solve the guiding problem

03:23 - Slew to AndXXI repeat OB1
  • put Red master and Blue slave
  • dofpia
  • stopped after 4th iteration - some astigmatism left

03:30 - AndXXI OB1 - repeat exposure
  • Red is master

03:47 - OB interrupted - the seeing is bad >1.5" (dimm 2.5")

03:48 - SWITCH TO LUCIFER
  • Pointing Correction by TO

04:16 - Slew to Campo A (Cresci)
  • Ks imaging 0001-0073
  • interrupted after 73 offsets, because it is complete

06:11 - Slew to Campo B (Cresci)
  • Ks imaging 0074-0148
  • Only 75 offsets executed
  • interrupted because airmass limit reached (airmass>1.5)
  • some elongation in the frames at low elevation

08:04 - Slew to PHOT_STD FS152
  • pointing correction
  • imaging in Ks 0149-153

08:26 - Slew to Pentericci Clustrer1
  • MOS acquisition 0153-0156

08:44 - Science spectroscopy script Cluster1 (Pentericci) 0157-0158
  • interrupted because it seemed out of focus
  • INITIALIZE
  • offset 0.0,1.5 (x,y) to bring the mask to initial position

08:59 - Repeat Science SpecScript for Clusted1 159-169
  • Sky lines are shifted and seem out of focus - INITIALIZE made no difference
  • Call Germany - they say they will run tests at the end of night
  • not clear if it moves between offsets or also during exposure (maybe that'a why it seems out of focus)
  • maybe something wrong with the flexure compensation
  • Script interrupted after 10 offsets + the 2 from before -> 12 offsets executed
  • OB is considered finished but it maybe compromised

09:53 - Calib Unit IN
  • ARC with Ar+Ne lamps 168-172
  • ARC with Ar+Xe lamps 0173-0176
  • FLAT 177-180

10:08 - Slew to SN2009kr image acquisition (Turatto) 0181-0182
  • Interrupted to put Flexure compensation OFF
For the rest of the night all imaging scripts have Flexure compensation OFF

10:19 - Repeat SN2009kr image acquisition (Turatto) 0183-0184

10:21 - Science imaging scripts for SN2009kr
  • Imaging H 0185-0200
  • Imaging J 0201-0208
  • Imaging script Ks 0209-0224

11:29 - Slew to SN2010sl for image acquisition 0225-0226
  • Science imaging script H 0227-0234
  • Science imaging script J 0235-0250
  • Science imaging script Ks 0251-0258
  • Sometimes it would give warnings about not reaching the guide star for the particular offset position

12:40 - Move to Flat position
  • used script to make the flats and they were saved as luci_20101011_Flat_)
  • Flats in J: 001-010 still dark, 011-020 low (4000), 031-040 high ( 18000-25000)
  • Flats in H: 021-030 low (6000-9000), 041-050 high ( 18000-23000)
  • Flats in Ks: 051-070 low (5000-8000), 071-090 high ( 14000-17000)

13:22 - END OF NIGHT
  • Hand over the telescopes to Germany for the tests to see what causes the shifting
  • They will run our Darks sciprt

-- VincenzoTesta - 11 Oct 2010
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback